Interest is compounded annually unless stated otherwise. Payments are at the end of the year unless stated otherwise. All bonds have a face value of $1000. Taxes are 0 unless specified otherwise.

3. Using market values, ABC has a debt ratio (debt/value) of .30. Equity has a return of 15% and debt is 4% and risk-free. The firm plans to issue equity to pay off debt. This will reduce the debt ratio to .15. The market risk premium is 10%. Find the return on equity after the capital structure change.
Answer #1

Calculating Beta,

15 = 4 + Beta(10)

Beta = 11/10

Beta = 1.1

Unlevered Beta = 1.1/(1 + 0.30)

Unlevered Beta = 0.846

Calculating new levered beta,

New Beta = 0.846(1 + 0.15)

New Beta = 0.9729

Return on Equity = 4 + 0.9729(10)

Return on Equity = 13.73%